### v1

@@ 2.2 What should we do next?
    Algorithms should be intellectually manageable: Although a common strategy in algorithm design is to first construct a program and then verify that it satisfies the required properties, this process is often cumbersome and error-prone. As Dijkstra (1972) famously quote:“ Program testing can be a very effective way to show the presence of bugs, but it is hopelessly inadequate for showing their absence. The only effective way to raise the confidence level of a program significantly is to give a convincing proof of its correctness. But one should not first make the program and then prove its correctness, because then the requirement of providing the proof would only increase the poor programmer’s burden.” Similarly, research on combinatorial methods—particularly BnB algorithms—often adopts the practice of constructing ad-hoc algorithms and proving their correctness post-hoc using complex proofs (or simply without proof). However, this approach is problematic. Complex algorithms typically result in complex proofs, which not only make the algorithms less accessible but also harder to verify. For instance, Demirović et al. (2022)’s algorithm spans nearly five pages of pseudocode, making it intellectually challenging to understand, implement, and verify. To address this, Dijkstra—and likewise, this work—advocates a constructive approach to programming: derive programs from well-defined specifications, thereby ensuring correctness by design. In combinatorial optimization, we begin with an unambiguous brute-force specification of the optimization problem and systematically derive a more efficient algorithm. This guarantees that the final implementation inherits the properties of the original specification.
@@
    Acceleration techniques must be demonstrably effective: This may seem counterintuitive: if a technique provably reduces the search space, how could it be ineffective? In practice, the overhead of applying the technique can outweigh its benefits. The key is to balance the cost of applying an acceleration method against the potential reduction in search space. For example, Brița et al. (2025) claim that caching is a major factor behind their algorithm’s efficiency. However, He and Little (2025b) provide a simple analysis showing that overlapping subproblems in decision tree construction are rare—particularly for the ODT problem with axis-parallel or more complex splitting rules—making caching or memoization largely ineffective. In our experiments, we even observed cases where Brița et al. (2025)’s algorithm explored over a million decision trees without a single cache hit, suggesting that in such scenarios, disabling caching entirely may improve performance.
@@
    Parallelism: For problems with intractable combinatorics—such as NP-hard optimization problems—parallelism is perhaps the only practical way to scale the algorithm to handle large scale datasetss. We do not merely advocate designing complex algorithms and parallelizing them post-hoc; rather, we argue that algorithms should be inherently parallelizable from the beginning of the design process. In other words, parallelizability should be treated as a fundamental property during algorithm design. Designing inherently parallel algorithms while maintaining optimality is challenging and requires a deep understanding of both the problem and algorithmic structure. The success of neural networks illustrates this well: GPU hardware significantly enhanced their effectiveness by aligning with their computational structure. Similarly, we argue that designing hardware-friendly algorithms is as important as achieving theoretical efficiency.
@@ 4.1 Definition of decision tree
Definition 1.
[…]
We call a decision tree that satisfies the following axioms, a proper decision tree with respect to $rs$:
@@
    Each branch node is defined by a single splitting rule $r:\mathcal{R}$, and each splitting rule subdivides the ambient space into two disjoint and connected subspaces, $r^{+}$ and $r^{-}$. The descendant rules of $r$ can be generated only from within $r^{+}$ and $r^{-}$,
@@
    Each leaf $L$ is defined by the intersection of subspaces $\bigcap_{p\in P_{L}}r_{p}^{\pm}$ for all the splitting rules $\left\{r_{p}\mid p\in P_{L}\right\}$ in the path $P_{L}:\mathcal{P}\left(\mathcal{A}\right)$ from the root to leaf $L$. The connected region (subspace) defined by $\bigcap_{p\in P_{L}}r_{p}^{\pm}$ is referred to as the decision region,
@@
    The ancestry relation between any pair of splitting rules $r_{i}\left(\swarrow\vee\searrow\right)r_{j}$ is transitive; in other words, if $r_{i}\left(\swarrow\vee\searrow\right)r_{j}$ and $r_{j}\left(\swarrow\vee\searrow\right)r_{k}$ then $r_{i}\left(\swarrow\vee\searrow\right)r_{k}$,
@@
    For any pair of splitting rules $r_{i}$ and $r_{j}$, only one of the following three cases is true: $r_{i}\swarrow r_{j}$, $r_{i}\searrow r_{j}$, and $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{j}$; additionally, $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{i}$ is always true;.
@@
Each hyperplane $h_{i}$ partitions $\mathbb{R}^{D}$ into two regions: the positive region $h_{i}^{+}=\left\{x\mid x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}\geq 0\right\}$ and the negative region $h_{i}^{-}=\left\{x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}<0\right\}$, where $\bar{x}=\left(x,1\right)\in\mathbb{R}^{D+1}$ denotes the data points in homogeneous coordinates. For brevity, we omit the labels on the branch arrows $\boldsymbol{w}^{T}x<0$ and $\boldsymbol{w}^{T}x\geq 0$, assuming by default that the left branch corresponds to the “less than” case and the right branch to the “greater than or equal to” case.
@@ 4.2.1 Defining combinatorial optimization problem formally through brute-force enumeration
In the theory of transformational programming (also known as constructive algorithmics) (Bird and De Moor, 1996), combinatorial optimization problems such as 1 is solved using the following, generic, generate-filter-select paradigm,
@@
A natural question arises: Can efficient solutions (efficient combinatorial optimization algorithms) always be derived from brute-force specifications, and how comprehensive is this derivation paradigm across combinatorial methods? Bird and De Moor (1996) formalized classical combinatorial methods, such as dynamic programming, greedy algorithms, and D&C, using a generic recursive program (hylomorphism) and derived efficient solutions as theorems. Building on this, He (2025) extended the framework to include BnB by incorporating search strategies into Bird’s theorems, further broadening the applicability of this approach. In practice, one only needs to verify the conditions of these theorems, after which program fusion follows automatically.
Furthermore, He (2025) established the following inclusion relationships among different combinatorial optimization methods:
@@
Where SDP stands for “sequential decision process,” a special form of recursion characterized by sequential and deterministic recursive steps. In simple, the inclusion relations in (7) shows that a general definition of D&C algorithms encompasses all classical combinatorial optimization methods. Both DP and classical D&C correspond to problems with overlapping and non-overlapping subproblems, respectively, and both can be viewed more generally than the BnB method, which is a sequential decision process combined with additional techniques, such as thinning and search strategies. Greedy algorithms represent the simplest instance, corresponding to the fused program of an SDP generator with the selector $\mathit{min}$. Importantly, this inclusion is not based on intuitive classification but on a rigorous formal abstraction of each method using abstract recursion, and these inclusion relations naturally emerge from the abstract level of these recursions. See Section II.2.9 of He (2025) for precise definitions of these terms and explanations.
@@ 4.2.2 Defining optimal decision tree problem through brute-force specifications
Proof.
See He and Little (2025b) for more details. ∎
@@
$\mathit{odt}_{\text{rec}}\left(K\right)$ $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{genDTKs}_{K}.$
@@
$\mathit{update}$ $:\mathcal{R}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\to\mathit{Maybe}\left(\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right)$
@@
Although the definition of e $\mathit{update}$ update appears complex, it is applied recursively along the path of the old tree $\mathit{DN}\left(u,s,v\right)$ to determine whether adding a new splitting rule $r$ would result in a non-proper tree. If the tree remains proper, rris added to the leaf; otherwise, $\mathit{Nothing}$ is returned, indicating a non-proper decision tree.
@@
One can check $\mathit{genDTs}_{\text{vec}}$ will generate the same set of proper decision trees as $\mathit{genDTs}_{\text{rec}}$, but will produce different list of trees. This is because different valid $K$-permutation can correspondent to the same tree (vice is not true). For instance, the tree
@@
Depth-constrained decision trees are often used in the study of ODT problems with smaller combinatorial complexity, such as axis-parallel decision trees (Mazumder et al., 2022; Brița et al., 2025) or decision trees over binary feature data Demirović et al. (2022); Verwer and Zhang (2019); Nijssen and Fromont (2007); Aglin et al. (2020); Nijssen and Fromont (2010). However, none of these studies formally define their problem or explain why their algorithm correctly solves the desired problem. In this subsection, we provide a formal characterization of decision trees defined in this manner.
@@
$\mathit{genDTDs}_{\text{depth}}$ $\left(d,\mathit{xs}\right)=\left[\mathit{DN}\left(u,r,v\right)\mid u\leftarrow\mathit{genDTDs}_{\text{depth}}\left(\mathit{xs}^{+},d-1\right),v\leftarrow\mathit{genDTDs}_{\text{depth}}\left(\mathit{xs}^{-},d-1\right),r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
This definition states that all possible decision trees of depth $d$ with respect to data list $\mathit{xs}$ can be constructed from decision trees of depth d1d-1 on the subsets $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$. Currently, a concrete definition for $\mathit{gen_{splits}}$ is not provided, as the splitting rule $r$ defining the decision tree problem has not yet been formalized. For the classical axis-parallel ODT problem, $\mathit{gen_{splits}}$ is straightforward: splitting rules can be generated from axis-parallel hyperplanes based on the data points. Equivalently, Brița et al. (2025); Mazumder et al. (2022) use the medians of adjacent pairs of data points for each dimension $D$.
@@
Some readers may find the recursion 23 closely related to the algorithm given by Brița et al. (2025), which has similar recursive structure. We will see shortly that Brița et al. (2025)’s algorithm can indeed be derived from the specification (24), which proves why Brița et al. (2025)’s algorithm is optimal.
4.2.3 Remarks of potential flaws for optimal decision tree problems over binary feature data
The recursions 23 and 18 are valid because, according to Axiom 1 of decision trees, subtrees can only be constructed from the subregions defined by $r$, which partition the data list $\mathit{xs}$ to $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$. This allows us to fix a root $r$, and safely divide the remaining splitting rules into two disjoint lists of splitting rules, they correspond to $\mathit{gen_{splits}}\left(\mathit{xs}^{\pm}\right)$ in $\mathit{odt}_{\text{depth}}$ and $r^{\pm}$ in $\mathit{odt}_{\text{rec}}$. In particular, we do not need to worry about elements of $\mathit{gen_{splits}}\left(\mathit{xs}^{\pm}\right)$ or $r^{\pm}$ appearing in $\mathit{gen_{splits}}\left(\mathit{xs}^{\mp}\right)$ or $r^{\mp}$ because of Axiom 4, which ensures that only one of the three possible cases occurs.
However, in the decision tree problem over binary feature data, a selected feature may appear in either the left or right subtree, depending on the presence or absence of the feature, regardless of the root $r$. Consequently, for any chosen root $r\in\mathit{gen_{splits}}\left(\mathit{xs}\right)$ (or $r\in\mathit{rs}$ for $\mathit{odt}_{\text{rec}}$) any subset of $\r\mathit{gen_{splits}}\left(\mathit{xs}\right)\backslash$ r or $\r\mathit{rs}\backslash$ r can form a feasible subtree. This substantially increases the complexity of splitting in binary-feature decision trees. It is therefore unclear why existing studies on ODT over binary feature data—such as Zhang et al. (2023); Demirović et al. (2022); Aglin et al. (2021); Lin et al. (2020); Hu et al. (2019); Verwer and Zhang (2019); Nijssen and Fromont (2010, 2007)—define splitting rules in ways that are either ambiguous or entirely unspecified.
For instance, Demirović et al. (2022) provide a relatively clear definition using a recursive formula (Equation (1) in their paper). Their method proceeds by sequentially scanning the features (denoted as $i\in\left[0,n-1\right]$ in Equation (1), where $n$ is the number of features) and partitioning the remaining elements by selecting a data point with index $i$. Here is our hypothesis: for each fix index $i$, Demirović et al. (2022) partition the index list $\left[0,n-1\right]$ into subsets of indices smaller than $i$ (size $i$) and greater than $i$ (size ni1n-i-1), reminiscent of Catalan-style splitting. However, the trees constructed over binary feature data do not satisfy the standard decision tree axioms: a descendant feature $j$ can simultaneously appear as both the left and right child of its ancestor $i$, and fixing a root $i$ does not uniquely determine how the remaining n1n-1 features are split. The only plausible explanation for their algorithmic process is that they implicitly fix an ordering of features by arranging them as a list $\left[0,n-1\right]$, which explains their use of list notation rather than set notation. In this sense, a fixed-order labeled set (i.e., a list) is equivalent to an unlabeled set, which justifies our description of their recursion as “Catalan-style splitting,” originally designed for unlabeled trees. Moreover, line 9 of Algorithm 2 in their paper explicitly uses a for-loop to scan all features, which further supports our interpretation. Therefore, unless Demirović et al. (2022) can demonstrate that such splitting rules exhaustively enumerate all possible labeled decision trees over binary feature data, their claim of optimality remains, at best, premature.
Also, Demirović et al. (2022) state that their algorithms can be considered starting points for their work, which may suggest that the same splitting technique was employed. Verwer and Zhang (2019) also adopt a splitting principle similar to that of Demirović et al. (2022).
Hu et al. (2019) define leaves as either unsplittable ($d_{\text{un}}$) or splittable ($d_{\text{split}}$), but the criteria for assigning leaves to $d_{\text{un}}$ remain unclear. Their recursive process first scans all possible combinations of $d_{\text{split}}$ and then scans subsets of these combinations. It is unclear why this two-step approach is necessary, as considering all subsets directly would be equivalent. Moreover, Hu et al. (2019) present a highly complex formula (Equation (1) in their paper) for counting the combinatorics of the decision tree problem over binary feature data, yet they provide neither explanation nor justification. Consequently, it remains unclear why the recursive process in their algorithm can exhaustively enumerate the recursion it describes. In their subsequent study, Lin et al. (2020) modify the splitting function to resemble the approach of Aglin et al. (2021); Nijssen and Fromont (2007, 2010) scanning all possible splitting rules sequentially and partitioning the remaining data using an undefined $\mathit{split}$ function (line 13-14 in Algorithm 1 GOSDT).
Without formal problem definitions and rigorous proofs of algorithmic optimality, these studies on ODT over binary feature data may be incorrect. Even if their final algorithms or implementations produce correct results, ambiguities in definitions can mislead newcomers and impede progress in the field. Therefore, it is highly desirable to provide formal proofs demonstrating the correctness—or incorrectness—of these algorithms. Unfortunately, this is currently impossible due to ambiguous problem definitions in these works

### v2

@@ 2.2 What should we do next?
    Algorithms should be intellectually manageable: For instance, Demirovic et al. (2022)’s algorithm spans nearly five pages of pseudocode, making it intellectually challenging to understand, implement, and verify. Although a common strategy in algorithm design is to first construct a program and then verify that it satisfies the required properties, this process is often cumbersome and error-prone. As Dijkstra (1972) famously quote:“ Program testing can be a very effective way to show the presence of bugs, but it is hopelessly inadequate for showing their absence. The only effective way to raise the confidence level of a program significantly is to give a convincing proof of its correctness. But one should not first make the program and then prove its correctness, because then the requirement of providing the proof would only increase the poor programmer’s burden.” Similarly, research on combinatorial methods—particularly BnB algorithms—often adopts the practice of constructing ad-hoc algorithms and proving their correctness post-hoc using complex proofs (or simply without proof). However, this approach is problematic. Complex algorithms typically result in complex proofs, which not only make the algorithms less accessible but also harder to verify To address this, Dijkstra—and likewise, this work—advocates a constructive approach to programming: derive programs from well-defined specifications, thereby ensuring correctness by design. In combinatorial optimization, we begin with an unambiguous brute-force specification of the optimization problem and systematically derive a more efficient algorithm. This guarantees that the final implementation inherits the properties of the original specification.
@@
    Acceleration techniques must be demonstrably effective: This may seem counterintuitive: if a technique provably reduces the search space, how could it be ineffective? In practice, the overhead of applying the technique can outweigh its benefits. The key is to balance the cost of applying an acceleration method against the potential reduction in search space. For example, Brita et al. (2025) claim that caching is a major factor behind their algorithm’s efficiency. However, He and Little (2025b) provide a simple analysis showing that overlapping subproblems in decision tree construction are rare—particularly for the ODT problem with axis-parallel or more complex splitting rules—making caching or memoization largely ineffective. In our experiments, we even observed cases where Brita et al. (2025)’s algorithm explored over a million decision trees without a single cache hit, suggesting that in such scenarios, disabling caching entirely may improve performance.
@@
    Parallelism: For problems with intractable combinatorics—such as NP-hard optimization problems—parallelism is perhaps the only practical way to scale the algorithm to handle large scale datasetss. We do not merely advocate designing complex algorithms and parallelizing them post-hoc; rather, we argue that algorithms should be inherently parallelizable from the beginning of the design process. In other words, parallelizability should be treated as a fundamental property during algorithm design. Designing inherently parallel algorithms while maintaining optimality is challenging and requires a deep understanding of both the problem and algorithmic structure. The success of neural networks illustrates this well: GPU hardware significantly enhanced their effectiveness by aligning with their computational structure. Similarly, we argue that designing hardware-friendly algorithms is as important as achieving theoretical efficiency. For example, in Part I, we show that solving the ODT problem with intractable splitting rules depends heavily on the algorithm’s ability to exploit parallelism. Furthermore, in Part II, we demonstrate that an algorithm with superior theoretical performance may still run slower than a less theoretically efficient one if the latter is better aligned with parallel hardware.
@@ 4.1 Definition of decision tree
Definition 1
Axioms for proper decision trees. We call a decision tree consists of splitting rules that satisfies the following axioms, a proper decision tree:
@@
    Structural constraint one (Ambient space partition): Each branch node is defined by a single splitting rule $r:\mathcal{R}$, and each splitting rule subdivides the ambient space into two disjoint and connected subspaces, $r^{+}$ and $r^{-}$.
@@
    Structural constraint two (Path intersection in leafs): Each leaf $L$ is defined by the intersection of subspaces $\bigcap_{p\in P_{L}}r_{p}^{\pm}$ for all the splitting rules $\left\{r_{p}\mid p\in P_{L}\right\}$ in the path $P_{L}$ from the root to leaf $L$. The connected region (subspace) defined by $\bigcap_{p\in P_{L}}r_{p}^{\pm}$ is referred to as the decision region.
@@
    Structural constraint three (Partition transitivity): Part I: The ancestry relation between any pair of splitting rules $r_{i}\left(\swarrow\vee\searrow\right)r_{j}$ is transitive; in other words, if $r_{i}\left(\swarrow\vee\searrow\right)r_{j}$ and $r_{j}\left(\swarrow\vee\searrow\right)r_{k}$ then $r_{i}\left(\swarrow\vee\searrow\right)r_{k}$. Part II: Moreover, $\boldsymbol{K}_{ij}=\pm 1$ if $r_{j}$ can be generated from $r_{i}^{\pm}$. As a result, any new decision rule $r$ added to a leaf must be generated within its corresponding decision region.
@@
    Ancestral constraint (Uniqueness of the ancestry relation): For any pair of splitting rules $r_{i}$ and $r_{j}$, only one of the following three cases is true: $r_{i}\swarrow r_{j}$, $r_{i}\searrow r_{j}$, and $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{j}$; additionally, $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{i}$ is always true; in other words, the possible value of $\boldsymbol{K}_{ij}\in\left\{1,0,-1\right\}$ is unique determined for all $i,j$ , and $\boldsymbol{K}_{ii}=0$ for all $i$.
A key observation in He and Little (2025b) is that if a tree’s splitting rules satisfy the structural constraints (Axioms 1-3), they define a class of decision tree problems. However, the ancestral constraint (Axiom 4), which specifies pairwise relationships between splitting rules, implicitly determines the subclass of decision tree problems under consideration. The proper decision tree framework adopts Axiom 4 as the core ancestral constraint because it encompasses most decision tree models in machine learning, including axis-parallel hyperplane, general hyperplane, and polynomial hypersurface decision trees. For decision trees over binary feature data, however, Subsection 4.4 explains how this problem violates Axiom 4, necessitating a different algorithmic approach. Fortunately, thanks to the deep algorithmic insights provided by our framework, this adaptation is achieved with minimal changes.
@@
Each hyperplane $h_{i}$ partitions $\mathbb{R}^{D}$ into two regions: the positive region
$h_{i}^{+}=\left\{x\mid x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}\geq 0\right\}$ and the negative region $h_{i}^{-}=\left\{x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}<0\right\}$, where $\bar{x}=\left(x,1\right)\in\mathbb{R}^{D+1}$ denotes the data points in homogeneous coordinates. For brevity, we omit the labels on the branch arrows $\boldsymbol{w}^{T}x<0$ and $\boldsymbol{w}^{T}x\geq 0$, assuming by default that the left branch corresponds to the “less than” case and the right branch to the “greater than or equal to” case.
@@ 4.2.1 Defining combinatorial optimization problem formally through brute-force enumeration
In the theory of transformational programming (also known as constructive algorithmics) (Bird and De Moor, 1996), combinatorial optimization problems such as (1) is solved using the following, generic, generate-filter-select paradigm,
@@
A natural question arises: Can efficient solutions (efficient combinatorial optimization algorithms) always be derived from brute-force specifications, and how comprehensive is this derivation paradigm across combinatorial methods? Bird and De Moor (1996) formalized classical combinatorial methods, such as dynamic programming, greedy algorithms, and D&C, using a generic recursive program (hylomorphism) and derived efficient solutions as theorems. Building on this, He (2025b) extended the framework to include BnB by incorporating search strategies into Bird’s theorems, further broadening the applicability of this approach. In practice, one only needs to verify the conditions of these theorems, after which program fusion follows automatically.
Furthermore, He (2025b) established the following inclusion relationships among different combinatorial optimization methods:
@@
where SDP stands for “sequential decision process,” a special form of recursion characterized by sequential and deterministic recursive steps. In simple, the inclusion relations in (7) shows that a general definition of D&C algorithms encompasses all classical combinatorial optimization methods. Both DP and classical D&C correspond to problems with overlapping and non-overlapping subproblems, respectively, and both can be viewed more generally than the BnB method, which is a sequential decision process combined with additional techniques, such as thinning and search strategies. Greedy algorithms represent the simplest instance, corresponding to the fused program of an SDP generator with the selector $\mathit{min}$.
Importantly, the inclusion (7) is not based on intuitive classification but on a rigorous formal abstraction of each method using abstract recursion, and these inclusion relations naturally emerge from the abstract level of these recursions. See Section II.2.9 of He (2025b) for precise definitions of these terms and explanations.
@@ 4.2.2 Defining optimal decision tree problem through brute-force specifications
Proof See He and Little (2025b) for more details.
@@
$\mathit{odt}_{\text{rec}}\left(K\right)$ $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{genDTSs}_{K}.$
@@
$\mathit{update}$ $:\mathcal{R}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\to\mathit{Maybe}\left(\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right)$ (21)
@@
Although the definition of e $\mathit{update}$ update appears complex, it is applied recursively along the path of the old tree $\mathit{DN}\left(u,s,v\right)$ to determine whether adding a new splitting rule $r$ would result in a non-proper tree. If the tree remains proper, $r$ is added to the leaf; otherwise, $\mathit{Nothing}$ is returned, indicating a non-proper decision tree.
@@
One can check $\mathit{genDTs}_{\text{vec}}$ will generate the same set of proper decision trees as $\mathit{genDTs}_{\text{rec}}$, but will produce different lists of trees. This is because different valid $K$-permutation can correspondent to the same tree (vice is not true). For instance, the tree
@@
Depth-constrained decision trees are often used in the study of ODT problems with smaller combinatorial complexity, such as axis-parallel decision trees (Mazumder et al., 2022; Brita et al., 2025) or decision trees over binary feature data Demirovic et al. (2022); Verwer and Zhang (2019); Nijssen and Fromont (2007); Aglin et al. (2020); Nijssen and Fromont (2010). However, none of these studies formally define their problem or explain why their algorithm correctly solves the desired problem. In this subsection, we provide a formal characterization of decision trees defined in this manner.
@@
$\mathit{genDTDs}_{\text{depth}}$ $\left(d,\mathit{xs}\right)=\big[\mathit{DN}\left(u,r,v\right)\mid$
$u\leftarrow$ $\mathit{genDTDs}_{\text{depth}}\left(\mathit{xs}^{+},d-1\right),v\leftarrow\mathit{genDTDs}_{\text{depth}}\left(\mathit{xs}^{-},d-1\right),r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\big]$
This definition states that all possible decision trees of depth $d$ with respect to data list $\mathit{xs}$ can be constructed from decision trees of depth d1d-1 on the subsets $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$. Currently, a concrete definition for $\mathit{gen_{splits}}$ is not provided, as the splitting rule $r$ defining the decision tree problem has not yet been formalized. For the classical axis-parallel ODT problem, $\mathit{gen_{splits}}$ is straightforward: splitting rules can be generated from axis-parallel hyperplanes based on the data points. Equivalently, Brita et al. (2025); Mazumder et al. (2022) use the medians of adjacent pairs of data points for each dimension $D$. A more detailed discussion of the various definitions of $\mathit{gen_{splits}}$ for solving different ML problems is provided in Subsection 5.2.
@@
Some readers may find the recursion 24 closely related to the algorithm given by Brita et al. (2025), which has similar recursive structure. We will see shortly that Brita et al. (2025)’s algorithm can indeed be derived from the specification (25), which proves why Brita et al. (2025)’s algorithm is optimal.